### 12 Dec 2023 v2

@@ 5.2 Not all cells observed
In this section we discuss the (common) case when not all cells are observed. We contribute three components. Firstly we detail how to use the reference validation approach in this case. Secondly, we propose a tool to validate the reference model. Thirdly, we propose a technique to combine the cell reference model score and LOCO cross validation model score.
@@ 5.2.1 Reference validation score
We begin by decomposing the population into two sets. The first set, $\mathbb{S}_{\mathrm{obs}}$, is the set of cells where at least one observation in each is observed in the sample. The total number of cells in this set is $S_{\mathrm{obs}}$ and any particular cell in this set is $s_{\mathrm{obs}}$. The second set, $\mathbb{S}_{\mathrm{unobs}}$, is the set of cells that do not have any observations in the sample. The total number of cells in this set is $S_{\mathrm{unobs}}$ and any particular cell in this set is $s_{\mathrm{unobs}}$. For clarity note that $J=S_{\mathrm{obs}}+S_{\mathrm{unobs}}$.
To compose the reference validation score, we note that when cells are observed we should use a LOCO approach to assess predictive power. When cells are not observed, we can simply use the difference in prediction between model $M_{c}$ and model $M_{*}$. This means we can write a score that combines PSIS approximated LOCO predictions with predictions for unobserved cells by modifying (29). For simplicity we denote
$\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x},w_{M_{c},j})=\frac{\sum_{b=1}^{B}w_{M_{c},j}^{b}\mbox{Pr}(Y_{i\in j}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin j},\textbf{x})}{\sum_{b=1}^{B}w_{M_{c},j}^{b}}$
@@
$\operatorname{Error}_{\mathrm{REF}}(M_{c},M_{*})^{2}=$
$\!\!\!\!\!\!\!\!\!\!\!\!\Bigg{(}\frac{1}{N}\Bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}\!\!N_{s_{\mathrm{obs}}}\bigg{(}\mbox{Pr}(Y_{i\in s_{\mathrm{obs}}}=1|\hat{\theta}^{b}_{M_{c}}y_{i\notin s_{\mathrm{obs}}},\textbf{x},w_{M_{c},s_{\mathrm{obs}}})-\mbox{Pr}(Y_{i\in s_{\mathrm{obs}}}=1|\hat{\theta}^{b}_{M_{*}}y_{i\notin s_{\mathrm{obs}}},\textbf{x},w_{M_{*},s_{\mathrm{obs}}})\bigg{)}+$
$\sum_{s_{unobs}=1}^{S_{unobs}}N_{s_{unobs}}\bigg{(}\mbox{Pr}(Y_{i\in s_{\mathrm{unobs}}}=1|\hat{\theta}_{M_{c}},y,\textbf{x})-\mbox{Pr}(Y_{i\in s_{\mathrm{unobs}}}=1|\hat{\theta}_{M_{*}},y,\textbf{x})\Bigg{)}\Bigg{)}^{2}.$ (30)
Each component of CRPS can similarly be decomposed into an approximate LOCO value for the cells with observations and a prediction the cells with no sample observations. We denote $g^{\prime\prime}_{\mathrm{REF}}(y_{j+1},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})=\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{\prime b},\textbf{x})$ and $h^{\prime\prime}_{\mathrm{REF}}(y,\textbf{x},\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{*}}^{b})=\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{c}}^{b},\textbf{x})-\mbox{Pr}(Y_{i\in j+1}|y,\hat{\theta}_{M_{*}}^{b},\textbf{x})$ as the corresponding $g(.)$ and $h(.)$ for the cells with no sample observations.
$\operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\ \ \ \ \frac{1}{B}\ \Bigg{(}\frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{c}}^{\prime b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})\bigg{)}\Bigg{|}\,+$
$\ \ \ \ \ \ \ \ +\ \frac{1}{2}\sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{RW,M_{*}}^{b},\hat{\theta}_{RW,M_{*}}^{\prime b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})\bigg{)}\Bigg{|}\,+$
$\ \ \ \ \ \ \ \ +\ \sum_{b=1}^{B}\Bigg{|}\frac{1}{N}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{obs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})+$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}N_{s_{\mathrm{unobs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{unobs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\bigg{)}\Bigg{|}\,\Bigg{)}.$ (31)
@@ 5.2.2 Combined validation score
Firstly we focus on the portion of the population that is observed. We subset to this section of the population and calculate the PSIS LOCO reference validation score and the PSIS LOCO cross validation score by treating the observed cells as a subpopulation; see Section 4. By comparing the resultant model ordering we can identify whether the two approaches suggest similar ordering for the subpopulation where the sample has an observation in each cell.
To demonstrate this, we extend the simulations in Section 2 to relax the assumption that every cell needs to be observed in the sample. Instead we simply constrain that every level of a variable needs to be observed in the sample when taking a sample from the population. With this constraint relaxed, our simulated samples contained $28.9\%$ of the cells in the population (min $24.9\%$ and max $33.1\%$). To begin with we focus on the portion of the population poststratification matrix observed in the sample.
Figure 8 plots the PSIS LOCO cross validation scores scores against the reference model scores. The first column uses the full model as the reference model, whilst the second column uses the precision model as the reference model. When the reference model is chosen appropriately (the full model) the two scores align, but when an inappropriate reference model is chosen (the precision model), they do not.
@@
However, this only evaluates part of the population and should not be used to score the models. Instead we propose an approach that combines the reference validation score and the cross validation score. In the reference validation approach we don’t leave out anything when we don’t have any observation in that cell, and in the combined approach the reference model approach is used only when there is no observation in that cell. For squared error we modify (30) to
$\operatorname{Error}_{\mathrm{COMB}}(M_{c},M_{*})^{2}=\Bigg{(}\frac{1}{N}\Bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}\bigg{(}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}\frac{\sum_{b=1}^{B}w_{s_{\mathrm{obs}}}^{b}\epsilon_{\mathrm{obs}}^{b}}{\sum_{b=1}^{B}w_{s_{\mathrm{obs}}}^{b}}\bigg{)}\ +$
$\ \ \ \ \sum_{s_{unobs}=1}^{S_{unobs}}\!\!N_{s_{unobs}}\!\bigg{(}\mbox{Pr}\left(Y_{i\in s_{\mathrm{unobs}}}=1\left|\hat{\theta}_{M_{c}},y,\textbf{x}\right.\right)-\mbox{Pr}\left(Y_{i\in s_{\mathrm{unobs}}}=1\left|\hat{\theta}_{M_{*}},y,\textbf{x}\right.\!\right)\!\Bigg{)}\!\Bigg{)}^{2}\!\!,$ (32)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \operatorname{CRPS}_{\mathrm{MRP},\mathrm{LOCO}}(M_{c},M_{*})=$
$\!\!\!\!\!\!\!\!\frac{1}{N}\Bigg{(}\sum_{b=1}^{B}\frac{1}{2B}\ \Bigg{|}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})+\!\!\!\!\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})\Bigg{|}\,+$
$\frac{1}{2}\ \Bigg{|}\sum_{s_{\mathrm{obs}}=1}^{S_{\mathrm{obs}}}N_{s_{\mathrm{obs}}}g(y_{i\notin s_{\mathrm{obs}}},\textbf{x},\hat{\theta}_{M_{*}}^{b},\hat{\theta}_{M_{*}}^{\prime b})+\!\!\!\!\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}g^{\prime\prime}_{\mathrm{REF}}(y_{s_{\mathrm{unobs}}},\textbf{x},y,\hat{\theta}_{M_{c}}^{b},\hat{\theta}_{M_{c}}^{\prime b})\Bigg{|}\,+$
$\Bigg{|}\sum_{s_{\mathrm{unobs}}=1}^{S_{\mathrm{unobs}}}\!\!N_{s_{\mathrm{unobs}}}h_{\mathrm{REF}}(y_{i\notin s_{\mathrm{unobs}}=1},\textbf{x},\hat{\theta}_{RW,M_{c}}^{b},\hat{\theta}_{RW,M_{*}}^{b})\Bigg{|}\,\Bigg{)}.$ (33)
We lose the observed cell component of the final term as there is no distribution for these estimates in the observed cell case.
@@ 6 Conclusion
In this manuscript we propose adaptions to the squared error and continuous ranked probability score (CRPS) for use in multilevel regression and poststratification (MRP) settings. We demonstrate that this adaption correctly recovers model ordering at the population and subpopulations. We also provide an approximate leave-one-cell-out approach for fast estimation. We propose a reference validation approach and a combined validation approach for use when not all cells are observed in the sample. Together we feel this provides encouraging progress towards MRP model validation and consider this a considerable improvement on previous model validation attempts (which do not retain correct ordering).
This work provides evidence that failure to correctly recover model ordering is based in the sum of individual scores. However, our method does consistently underestimate the magnitude of the loss, which should be further investigated. Until a solution has been identified, the magnitude of loss should not be directly used as an approximation of loss of the true error.
In this manuscript we have focused primarily on the simulation design proposed by (Kuh et al., 2022). However, future work should consider different sampling conditions and model comparisons to ensure generalisability. This example is severe and may not be illustrative of real world applications. It could be useful to consider the scoring of models with different priors.

### 12 Apr 2024 v3

@@ 5.2 Not all cells observed
Previously, we demonstrated the equivalence of a reference validation approach to a cross validation approach when all cells are observed. Here, we relax the constraint that all cells are observed. We contribute three components. Firstly we detail how to use the reference validation approach in this case. Secondly, we propose a technique to validate the reference model. Thirdly, we propose a combined validation method using LOCO cross validation and LOCO reference validation.
@@ 5.2.1 Reference validation score
We begin by partitioning the population into two sets. The first set, $\mathbb{V}$, is the set of cells where at least one observation in each is observed in the sample. The total number of cells in this set is $V$ and any particular cell in this set is $v$. The second set, $U$, is the set of cells that do not have any observations in the sample. The total number of cells in this set is $U$ and any particular cell in this set is $u$. For clarity note that $J=V+U$.
When cells are observed, we could use a cross validation or reference validation approach to assess predictive power. When cells are not observed, we can only use a reference validation approach. To minimise the reliance on correctly specifying the reference model, we could combine cross validation for observed cells and reference validation for unobserved cells. For simplicity, we write the Pareto smoothed LOCO approximation as
$\mbox{Pr}(Y_{i\in j}|\hat{\theta}^{b}_{c},w_{c,j})=\frac{\sum_{b=1}^{B}w_{c,j}^{b}\mbox{Pr}(Y_{i\in j}|\hat{\theta}^{b}_{c})}{\sum_{b=1}^{B}w_{c,j}^{b}}.$
@@
$\operatorname{Error}^{(\mathrm{REF})}(M_{c},M_{*})=\frac{1}{N}\Bigg{(}\sum_{v=1}^{V}\!\!N_{v}\bigg{(}\mbox{Pr}(Y_{i\in v}|\hat{\theta}^{b}_{c},w_{c,v})-\mbox{Pr}(Y_{i\in v}|\hat{\theta}^{b}_{*},w_{*,v})\bigg{)}+\\ \sum_{u=1}^{V}N_{u}\bigg{(}\mbox{Pr}(Y_{i\in u}|\hat{\theta}_{c})-\mbox{Pr}(Y_{i\in u}|\hat{\theta}_{*})\Bigg{)}.$ (23)
Each component of CRPS can similarly be decomposed into an approximate LOCO reference validation for the cells with observations and approximate LOCO reference validation for the cells with no sample observations. We denote $g_{j}^{(\mathrm{REF})}(.)$ and $h_{j}^{(\mathrm{REF})}(.)$ as the corresponding $g_{j}(.)$ and $h_{j}(.)$ for the reference model approach.
$\widehat{\operatorname{CRPS}}^{(\mathrm{REF})}(M_{c},M_{*})=\\ \frac{1}{BN}\sum_{b=1}^{B}\Bigg{(}\frac{1}{2}\Bigg{|}\sum_{v=1}^{V}N_{v}g_{v}^{(\mathrm{REF})}(\hat{\theta}_{c}^{b},\hat{\theta}_{c}^{\prime b})+\sum_{u=1}^{U}N_{u}g_{u}^{(\mathrm{REF})}(\hat{\theta}_{c}^{b},\hat{\theta}_{c}^{\prime b})\Bigg{|}\,+\\ +\ \frac{1}{2}\Bigg{|}\sum_{v=1}^{V}N_{v}g_{v}^{(\mathrm{REF})}(\hat{\theta}_{*}^{b},\hat{\theta}_{*}^{\prime b})+\sum_{u=1}^{U}N_{u}g_{u}^{(\mathrm{REF})}(\hat{\theta}_{*}^{b},\hat{\theta}_{*}^{\prime b})\Bigg{|}\,-\\ \ \Bigg{|}\sum_{v=1}^{V}N_{v}h_{v}^{(\mathrm{REF})}(\hat{\theta}_{c}^{b},\hat{\theta}_{*}^{b})+\sum_{u=1}^{U}N_{u}h_{u}^{(\mathrm{REF})}(\hat{\theta}_{c}^{b},\hat{\theta}_{*}^{b})\Bigg{|}\,\Bigg{)}.$ (24)
@@ 5.2.2 Combined validation score
We begin by focusing on the portion of the population that is observed. We calculate the PSIS LOCO reference validation score and the PSIS LOCO cross validation score by treating the observed cells as a subpopulation; see Section 4. By comparing the resultant model ordering we can identify whether the two approaches suggest similar ordering for the subpopulation where the sample has an observation in each cell.
To demonstrate this, we extend the simulations in Section 2 to relax the assumption that every cell needs to be observed in the sample. Instead we simply constrain that every level of a variable needs to be observed in the sample when taking a sample from the population. With this constraint relaxed, our simulated samples contained $29\%$ of the cells in the population (minimum $25\%$ and maximum $33\%$).
Figure 5 plots the PSIS LOCO cross validation scores scores against the reference model scores. The first column uses the full model as the reference model, whilst the second column uses the precision model as the reference model. When the reference model is chosen appropriately (the full model) the ordering of scores align, but when an inappropriate reference model is chosen (the precision model), they do not.
@@
However, this only evaluates part of the population and should not be used to score the models. Instead we propose an approach that combines the reference validation score and the cross validation score. We use the reference validation approach when we don’t have an observation in a cell, and the cross validation approach when we do. For squared error we modify (23) to
$\widehat{\operatorname{Error}}^{(\mathrm{CB})}(M_{c},M_{*})=\frac{1}{N}\bigg{(}\sum_{v=1}^{V}N_{v}\frac{\sum_{b=1}^{B}w_{v}^{b}\epsilon_{v}^{b}}{\sum_{b=1}^{B}w_{v}^{b}}\ +\sum_{u=1}^{U}\!\!N_{u}\!\bigg{(}\mbox{Pr}(Y_{i\in u}|\hat{\theta}_{c})-\mbox{Pr}(Y_{i\in u}|\hat{\theta}_{*})\bigg{)}\!\!\bigg{)}\!\!,$
@@
$\widehat{\operatorname{CRPS}}^{(\mathrm{CB})}(M_{c},M_{*})=\\ \frac{1}{BN}\sum_{b=1}^{B}\Bigg{(}\frac{1}{2}\Bigg{|}\sum_{v=1}^{V}N_{v}g_{v}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{c,RW}^{\prime b})+\sum_{u=1}^{U}N_{u}g_{u}^{(\mathrm{REF})}(\hat{\theta}_{c,RW}^{b},\hat{\theta}_{c,RW}^{\prime b})\Bigg{|}\,+\\ +\ \frac{1}{2}\Bigg{|}\sum_{v=1}^{V}N_{v}g_{v}(\hat{\theta}_{*,RW}^{b},\hat{\theta}_{*,RW}^{\prime b})+\sum_{u=1}^{U}N_{u}g_{u}^{(\mathrm{REF})}(\hat{\theta}_{*}^{b},\hat{\theta}_{*}^{\prime b})\Bigg{|}\,-\left|\sum_{u=1}^{U}N_{u}h_{u}^{(\mathrm{REF})}(\hat{\theta}_{c}^{b},\hat{\theta}_{*}^{b})\right|\Bigg{)}.$ (25)
There is one less term on the final line as there is no distribution for these estimates in the observed cell case.
@@ 6 Conclusion
We have consider the challenge of assessing model adequacy and selection in the case where models are being used to predict and then aggregate. Our specific focus is on multilevel regression and poststratification (MRP), but this challenge is relatively common across a range of fields. We propose adaptions to the squared error and continuous ranked probability score (CRPS) and demonstrate that this adaption correctly recovers model ordering at the population and subpopulations, suggesting a successful tool for selecting models. We also provide an approximate leave-one-cell-out approach for fast estimation. We propose a reference validation approach and a combined validation approach for use when not all cells are observed in the sample. Together we feel this provides encouraging progress towards MRP model validation and consider this a considerable improvement on previous model validation attempts (which do not retain correct ordering).
This work provides evidence that failure to correctly recover model ordering is based in the sum of individual scores. However, our method does consistently underestimate the magnitude of the loss, which should be further investigated. Until a solution has been identified, the magnitude of loss should not be directly used as an approximation of true loss, suggesting that caution should be used when evaluating model adequacy. This work used the simulation design proposed by Kuh et al. (2023). Future work could consider different sampling conditions and model comparisons (including comparisons of different priors) to ensure generalisability as this scenario is severe and may not be illustrative of real world applications.